OpenStudy (anonymous):

Suppose you mark n points on a circle, where n is a whole number greater than 1. The number of segments you can draw that connect these points is 1/2n2 - 1/2n How many segments can you draw if you mark 8 points on the circle?

OpenStudy (anonymous):

i assume this is \[\frac{1}{2}n^2-\frac{1}{2}n\] replace \(n\) by \(8\) and compute \[\frac{1}{2}\times 8^2-\frac{1}{2}\times 8\]

OpenStudy (anonymous):

is it 32 segments

OpenStudy (radar):

First term = 32, but you are not completed, there is a second term.